# Adaptive NAD: Online and Self-adaptive Unsupervised Network Anomaly Detector

### v2

@@ Abstract
The widespread usage of the Internet of Things (IoT) has raised the risks of cyber threats, thus developing Anomaly Detection Systems (ADSs) that can adapt to evolving or new attacks is critical. Previous studies primarily focused on offline unsupervised learning methods to safeguard ADSs, which is not applicable in practical real-world applications. Besides, most of them strongly rely on assumptions of known legitimates and fail to satisfy the interpretable requirements in security applications, creating barriers to the adoption in practice. In this paper, we design Adaptive NAD, a general framework to improve and interpret online unsupervised anomaly detection in security domains. An interpretable two-layer anomaly detection strategy is proposed to generate reliable high-confidence pseudo-labels. Then, an online learning scheme is introduced to update Adaptive NAD by a novel threshold calculation technique to adapt to new threats. Experimental results demonstrate that Adaptive NAD achieves more than 5.4%, 23.0%, and 3.2% improvements in SPAUC compared with state-of-the-art solutions on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. The code is released at https://github.com/MyLearnCodeSpace/Adaptive-NAD.
@@ I Introduction
Primarily, existing unsupervised anomaly detection approaches strongly rely on assumptions of known legitimate [8, 9, 10, 11, 12, 13, 14, 11, 15, 16, 17] and/or even malicious data [18, 19] to support the training the initial unsupervised model. Moreover, recent existing unsupervised-learning-based research [8, 9, 10, 20, 15] conducts anomaly detection by using a fixed threshold, which tends to fail to detect anomalies when concept drifts exist. To address this, a dynamic threshold is utilized to distinguish normal from anomalies in [11, 21], but suffers from high false positive rates. Additionally, most existing literature [8, 9, 10, 12, 14, 15, 22, 16] trains their unsupervised models offline without any adaption to drifting patterns, making them not applicable in real-world applications. There is only limited research on online unsupervised anomaly detection, and most of them like [23, 24, 25] consider “online” as the model being trained offline and utilized for real-time inferences without model updating.
@@
Our work. To address the above issues, we propose a general framework, named Adaptive NAD, to interpret and improve online unsupervised deep learning-based anomaly detection in security applications. The high-level goal of Adaptive NAD is to design a novel interpretable online unsupervised framework for anomaly detection that meets several practical security requirements (such as generalization, human-readable, low false positives, and adapting new threats). To this end, we designed two key strategies in Adaptive NAD: (i) interpretable two-layer based unsupervised anomaly detection strategy leveraging advantages of both supervised and unsupervised models to generate highly confident pseudo labels in highly imbalanced network streaming data. (ii) Online learning strategy for fine-tuning models and regenerating loss distributions aligned with the evolving network events to alleviate the effect of concept drift.
@@ III-A Interpretable Two-layer Anomaly Detection Strategy (ITAD-S)
High-confidence pseudo label generation process: This process utilizes a novel Adaptive Threshold Calculation technique to select high-confidence labeled samples predicted by the unsupervised learning model from the input data. It is utilized to eliminate manual labeling for autonomous updates. Our framework does not assume any known normal or anomalous data. Given the fact that anomalies are usually much less than normal samples in practical security-related applications. Unsupervised learning models together with the Adaptive Threshold Computation technique are used to generate the high confidence pseudo-labels.
@@ III-C Adaptive Threshold Calculation Technique
When new network samples arrive, the behavioral patterns of network traffic may change. Consequently, the predefined thresholds for classifying network samples (e.g., normal or abnormal) need to be periodically updated. Therefore, a novel dynamic threshold calculation strategy that automatically learns from the ongoing behaviors and adapts the corresponding thresholds to newer coming data. The initial thresholds $T_{1}$ and $T_{2}$ are calculated based on a limited dataset (e.g., the data from the first day or randomly selected 1% from the dataset). Given that the ground-truth labels of the first-day dataset are unknown, we assume all samples are normal (this assumption is reasonable since, in anomaly detection systems, the vast majority of samples are normal). We use these “normal” samples to train the unsupervised learning model. Initially, because we assume all samples are normal, we use only one threshold $T_{1}$. Then, in the second round of training, we use the trained unsupervised model to predict the loss values for new samples. These loss values are divided into normal and abnormal samples using the threshold $T_{1}$. When a sufficient number of abnormal samples are collected (e.g., when the number of abnormal samples reaches 500), we calculate a second threshold $T_{2}$. From the third round of training onward, we follow the process described in Section II for training and updating. In this way, the threshold for the unsupervised model $M$ can be automatically determined for ongoing network anomaly detection.
@@ IV-A Dataset
The Edge-IIoTset dataset [29] is designed to evaluate intrusion detection systems (IDS) in IoT and IIoT environments. It is generated from a purpose-built testbed featuring diverse devices, sensors, protocols, and cloud/edge configurations. The dataset includes fourteen attack types related to IoT and IIoT connectivity protocols. his paper evaluates Adaptive NAD on the benign traffic and attack traffic binary classification task. which has 2,154,864 benign and 1,076,200 attack entities. The dataset was divided into three subsets: a first-round training set (1% of the total data, with 21,548 benign and 10,762 attack data), a training set (69% of the total data, with 1,486,856 benign and 742,578 attack data), and a testing set (30% of the total data, with 646,460 benign and 322,860 attack data).
@@ IV-B Setup
System settings. Our NAD setup includes an encoder and a decoder, where both the encoder and the decoder have one LSTM layer with 64 hidden neurons, and the decoder has an additional fully connected (FC) layer. In the CIC-Darknet2020 [27], our timestep is 30 and update the model every 6,400 samples, while in both the CIC-DoHBrw-2020 [28] and Edge-IIoTset [29], our timestep is 5 and update the model in every 64,000 samples. The input to NAD is network traffic data, and the latent dimension is 32. All simulations are carried out on a Linux server, with AMD EPYC 7282 16-Core Processor at 2.80 GHz and 377 GB of RAM. We implement the ADS using the Adaptive NAD framework in Python and construct the models with Pytorch and Scikit Learn libraries. All algorithms are implemented in the same environment, and the average results of 10 runs are collected for fair comparison.
@@ V Discussions
Limitations and future work. Firstly, LSTM-VAE and RF are used in the current framework. More models could be attempted in the future to select as first or second layer of Adaptive NAD for optimizing the performance of the proposed framework. Secondly, we configure the hyper-parameters of Adaptive NAD empirically, which lacks systematic strategies to configure them. Future work can investigate more techniques to configure the hyper-parameters. Thirdly, two public datasets are employed to verify the proposed model. In the future, we intend to gather and integrate network events from different real-world systems and use these data to further evaluate the robustness of ADSs built using the Adaptive NAD framework. Fourthly, in addition to the applications in the security domain, our work could be generalized to other domains for detecting outliers.

### v3

@@ Abstract
The widespread usage of the Internet of Things (IoT) has raised the risks of cyber threats, thus developing Anomaly Detection Systems (ADSs) that can adapt to evolving or new attacks is critical. Previous studies primarily focused on offline unsupervised learning methods to safeguard ADSs, which is not applicable in practical real-world applications. Besides, most of them strongly rely on assumptions of known legitimates and fail to satisfy the interpretable requirements in security applications, creating barriers to the adoption in practice. In this paper, we design Adaptive NAD, a general framework to improve and interpret online and Adaptive unsupervised Network Anomaly Detection in security domains. An interpretable two-layer anomaly detection strategy is proposed to generate reliable high-confidence pseudo-labels. Then, an online learning scheme is introduced to update Adaptive NAD by a novel threshold calculation technique to adapt to new threats. Experimental results demonstrate that Adaptive NAD achieves more than 5.4%, 23.0%, and 3.2% improvements in SPAUC compared with state-of-the-art solutions on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. The code is released at https://github.com/MyLearnCodeSpace/Adaptive-NAD.
@@ I Introduction
Primarily, existing unsupervised anomaly detection approaches strongly rely on assumptions of known legitimate [8, 9, 10, 11, 12, 13, 14, 11, 15, 16, 17] and/or even malicious data [18, 19] to support the training of the initial unsupervised model. Moreover, recent existing unsupervised-learning-based research [8, 9, 10, 20, 15] conducts anomaly detection by using a fixed threshold, which tends to fail to detect anomalies when concept drifts exist. To address this, a dynamic threshold is utilized to distinguish normal from anomalies in [11, 21], but suffers from high false positive rates. Additionally, most of existing literature [8, 9, 10, 12, 14, 15, 22, 16] trains their unsupervised models offline without any adaption to drifting patterns, making them not applicable in real-world applications. There is only limited research on online unsupervised anomaly detection, and most of them like [23, 24, 25] consider “online” as the model being trained offline and utilized for real-time inferences without model updating.
@@
Our work. To address the above issues, we propose a general framework, named Adaptive NAD, to interpret and improve online unsupervised deep learning-based anomaly detection in security applications. The high-level goal of Adaptive NAD is to design a novel interpretable online unsupervised framework for anomaly detection that meets several practical security requirements (such as generalization, human-readable, low false positives, and adapting new threats). To this end, we designed two key strategies in Adaptive NAD: (i) interpretable two-layer based unsupervised anomaly detection strategy by leveraging advantages of both supervised and unsupervised models to generate highly confident pseudo labels in highly imbalanced network streaming data. (ii) Online learning strategy for fine-tuning models and regenerating loss distributions aligned with the evolving network events to alleviate the effect of concept drift.
@@ III-A Interpretable Two-layer Anomaly Detection Strategy (ITAD-S)
High-confidence pseudo label generation process: This process utilizes a novel Adaptive Threshold Calculation technique to select high-confidence labeled samples predicted by the unsupervised learning model from the input data. It is utilized to eliminate manual labeling for autonomous updates. Our framework does not assume any known normal or anomalous data. Given the fact that anomalies are usually much less than normal samples in practical security-related applications.
@@ III-C Adaptive Threshold Calculation Technique
When new network samples arrive, the behavioral patterns of network traffic may change. Consequently, the predefined thresholds for classifying network samples (e.g., normal or abnormal) need to be periodically updated. Therefore, a novel dynamic threshold calculation strategy that automatically learns from the ongoing behaviors and adapts the corresponding thresholds to newer coming data. The initial thresholds $T_{1}$ and $T_{2}$ are calculated based on a limited dataset (e.g., the data from the first day or randomly selected 1% from the dataset). We use these “normal” samples to train the unsupervised learning model. Initially, because we assume all samples are normal, we use only one threshold $T_{1}$. Then, in the second round of training, we use the trained unsupervised model to predict the loss values for new samples. These loss values are divided into normal and abnormal samples using the threshold $T_{1}$. When a sufficient number of abnormal samples are collected (e.g., when the number of abnormal samples reaches 500), we calculate a second threshold $T_{2}$. From the third round of training onward, we follow the process described in Section II for training and updating. In this way, the threshold for the unsupervised model $M$ can be automatically determined for ongoing network anomaly detection.
@@ IV-A Dataset
The Edge-IIoTset dataset [29] is designed to evaluate intrusion detection systems (IDS) in IoT and IIoT environments. It is generated from a purpose-built testbed featuring diverse devices, sensors, protocols, and cloud/edge configurations. The dataset includes fourteen attack types related to IoT and IIoT connectivity protocols. This paper evaluates Adaptive NAD on the benign traffic and attack traffic binary classification task. which has 2,154,864 benign and 1,076,200 attack entities. The dataset was divided into three subsets: a first-round training set (1% of the total data, with 21,548 benign and 10,762 attack data), a training set (69% of the total data, with 1,486,856 benign and 742,578 attack data), and a testing set (30% of the total data, with 646,460 benign and 322,860 attack data).
@@ IV-B Setup
System settings. Our NAD setup includes an encoder and a decoder, where both the encoder and the decoder have one LSTM layer with 64 hidden neurons, and the decoder has an additional fully connected (FC) layer. We use the Adam optimizer with a learning rate of 0.001. In the CIC-Darknet2020 [27], our timestep is 30 and update the model every 6,400 samples, while in both the CIC-DoHBrw-2020 [28] and Edge-IIoTset [29], our timestep is 5 and update the model in every 64,000 samples, which are the optimal configuration determined through extensive experiments. The input to NAD is network traffic data, and the latent dimension is 32. Based on our preliminary experiments, we calculated $T_{1}$ of the CIC-Darknet2020 [27] dataset using the 98th percentile and $T_{2}$ using the 10th percentile. For the CIC-DoHBrw-2020 [28] and Edge-IIoTset [29] datasets, we selected the 95th percentile to calculate $T_{1}$ and the 1st percentile to calculate $T_{2}$.
All simulations are carried out on a Linux server, with AMD EPYC 7282 16-Core Processor at 2.80 GHz and 377 GB of RAM. We implement the ADS using the Adaptive NAD framework in Python and construct the models with Pytorch and Scikit Learn libraries. All algorithms are implemented in the same environment, and the average results of 10 runs are collected for fair comparison.
@@ V Discussions
Limitations and future work. Firstly, LSTM-VAE and RF are used in the current framework. More models could be attempted in the future to select as first or second layer of Adaptive NAD for optimizing the performance of the proposed framework. Secondly, we configure the hyper-parameters of Adaptive NAD empirically, which lacks systematic strategies to configure them. Future work can investigate more techniques to configure the hyper-parameters. Thirdly, three public datasets are employed to verify the proposed model. In the future, we intend to gather and integrate network events from different real-world systems and use these data to further evaluate the robustness of ADSs built using the Adaptive NAD framework. Fourthly, in addition to the applications in the security domain, our work could be generalized to other domains for detecting outliers.